A Chinese problem states that several people combined their money to buy a farm tool to share. If each person paid 8 coins, they had 3 too many. If each person paid 7 coins, there were 4 coins too short. How many people were there, and what was the price of the farm tool?
Answer by macston(5194)

.
y=cost of tool; x=number of people
.
y=8x-3
y=7x+4
.
8x-3=7x+4
x-3=4
x=7
.
ANSWER 1: There were 7 people.
.
y = 8x-3 = 8(7)-3 = 56-3 = 53
.
ANSWER 2: The tool cost 53 coins.
.
CHECK:
y=7x+4
53=7(7)+4
53=49+4
53=53
.